# Additive subtraction games

### 11 Mar 2026 v1

@@ Abstract.
We determine the full nim-value structure of additive subtraction games in the primitive quadratic regime. The problem appears in Winning Ways by Berlekamp et al. in 1982; it includes a closed formula, involving Beatty-type bracket expressions on rational moduli, for determining the $\mathscr{P}$-positions, but to the best of our knowledge, a complete proof of this claim has not yet appeared in the literature; Miklós and Post (2024) established outcome-periodicity, but without reference to that closed formula. The primitive quadratic case captures the source of the quadratic complexity of the problem, a claim supported by recent research in the dual setting of sink subtraction with Bhagat et al. This study focuses on a number theoretic solution involving the classical closed formula, and we establish that each nim-value sequence resides on a linear shift of the classical $\mathscr{P}$-positions.
@@ 1. Introduction
Subtraction games are two-player games, played on heaps of counters, and players alternately remove a number $s$ of counters, provided $s\in S$, a given set of positive integers that remain constant throughout play. The positions are encoded by the natural numbers and a player who cannot move, because every subtraction leads to a negative number, loses. We study the nim-values of additive subtraction. Let $S=\{a,b,a+b\}$ with $b=a+\delta$ and assume the primitive quadratic regime where $a<\delta<2a$ and $\gcd(a,\delta)=1$. The nim-values (a.k.a nimbers), are game theoretic natural numbers recursively defined as the smallest non-negative value that does not appear among the options of a given game. For example, if the options have values in the set $\{0,1,3\}$, then the game position has nim-value two. This is the classical mex rule, and we think of it in terms of its two components: given a candidate set of nim-valued positions, for each position, one must verify both anti-collision (the same value does not appear among the options) and reachability (every smaller value appears among the options). The zero-valued positions are also known as the $\mathscr{P}$-positions.
A classical closed form (see e.g. [2, 5]) simplifies in this primitive quadratic regime to the following claimed $\mathscr{P}$-positions: for all $n\in\mathbb{N}_{0}$, let
(1) $w_{n}:=n+a\Big\lfloor\frac{n}{a}\Big\rfloor+b\Big\lfloor\frac{n}{\delta}\Big\rfloor.$
To the authors’ best knowledge, no proof of this claim has yet appeared in the literature (the formula first appears in [2] but without a proof). For this candidate set of $\mathscr{P}$-positions, we write
$\mathcal{W}_{0}=\{w_{n}:n\geqslant 0\}.$
We call an expression of the form
(2) $\sum_{i=1}^{t}c_{i}\left\lfloor n\alpha_{i}\right\rfloor,$
with integer coefficients $c_{i}$ and positive rational moduli $\alpha_{i}$, a bracket expression on $t$ terms. Note that in (1), the leading term $n=1\cdot\lfloor n/1\rfloor$ is itself a bracket term with $c_{1}=\alpha_{1}=1$. So this is a bracket expression on three terms.
In Section 2, we begin by justifying anti-collision: there is no move between any two positions in $\mathcal{W}_{0}$. Later we will explain the full characterization of all nim-values in $\{0,1,2,3\}$; since the ruleset has three move options the mex-rule implies that there can be no larger nim-values. It turns out that the non-zero nim-values are affine translations of $\mathcal{W}_{0}$, but with an interesting restriction in the case of nim-value three positions. We require elementary number theory to count the number of such positions, and we devote a separate section to this, Section 4. Positions of nim-value two are also devoted to a separate section, Section 3, while nim-value one positions are automatic by using Ferguson’s pairing principle, which was already mentioned in [2].
Let us remark that “periodicity” can be measured via an index-period (counting indexes of a sequence) or a heap-period (counting game positions). As we will see, the arithmetic of the problem ultimately stems from the regular gap structure of the $\mathscr{P}$-position sequence (1): only four gap sizes appear, and their locations are determined exactly by the residue classes of the index modulo $a$ and $\delta$. It is easy to see that the index-period of (1) is $a\delta$, while the heap-period is $(3\delta+a)a$; see also (10).
The nim-value classification is established by a single strong induction on the position size: assuming that for all position smaller than $x$ the nim-values have been correctly assigned, we verify the mex rule consistency at position $x$. In Section 6 we assemble all ingredients into the proof of the following result. If $X$ is a set of natural numbers and $y$ is a natural number, then $X+y:=\{x+y:x\in X\}$.
@@
Assume $a<\delta<2a$ and $\gcd(a,\delta)=1$, with $b=a+\delta$. Then $\mathcal{W}_{0}$ is the set of $\mathscr{P}$-positions of the additive subtraction game $\{a,b,a+b\}$, and the set of nim-value one positions is $\mathcal{W}_{0}+a$. The set of nim-value two positions is $\mathcal{W}_{0}-b$, and the set of nim-value three positions is $(\mathcal{W}_{0}-\delta)\setminus\mathcal{W}_{0}$. Together, these sets partition the natural numbers.
The primitive quadratic regime, where $a<\delta<2a$, is the first parameter range in which the two scales $a$ and $\delta$ interact non-trivially in the bracket expression. Outside this regime, essentially nothing new occurs: either one denominator becomes a multiple of the other, reducing the problem to linear complexity, or certain homogeneous regions of nimbers expand regularly by $2a$, preserving quadratic complexity. We omit further discussion of those cases here. The work of Bhagat et al. [3] in the dual sink subtraction setting further supports the view that the primitive quadratic regime captures the core complexity of the general theory.
@@ 2. Anti-collision of candidate $\mathscr{P}$-positions
In this section, we prove the first part of the mex-verification for a candidate set of $\mathscr{P}$-positions, anti-collision. Note that the reachability part is empty, since we here study zero values. However, since any independent subset of the natural numbers satisfies the anti-collision requirement, the true positioning of the zero values will only be confirmed when we have established the complete nim-value tetra-partitioning of all game positions, as in Theorem 1. Concerning the induction proof, note that the nim-value of position $w_{0}=0$ is zero.
@@
Let $S=\{a,b,a+b\}$ with $b=a+\delta$ and $a<\delta<2a$. For all $n\in\mathbb{N}_{0}$, let
@@
If $\beta\geqslant 1$ then $b(\beta-1)\geqslant 0$, and since $a\alpha\geqslant 0$ and $\ell\geqslant 1$ we reach a contradiction. Hence $\beta=0$, which gives
@@
Let us define $\mathcal{W}_{1}=\{w_{n}+a:n\geqslant 0\}$. Every position $x$ in this set has an option in $\mathcal{W}_{0}$, namely xax-a. If the set $\mathcal{W}_{0}$ constitutes the set of $\mathscr{P}$-positions, then, since $a=\min S$, Ferguson’s pairing principle [2] implies that every position in $\mathcal{W}_{1}$ has nim-value one, and there is no other position of value one.
@@ 3. Higher nim-values
In this section, we study positions with nim-values two and three. Define, for all $n\in\mathbb{N}_{0}$, $v_{n}:=w_{n}-b$ and let $\mathcal{W}_{2}:=\{v_{n}:n\geqslant 0\}$. The following lemma identifies the key arithmetic property of positions in $\mathcal{W}_{2}$. It makes explicit the effect of subtracting $b$ from any element of $\mathcal{W}_{2}$; observe that $v_{n}-b=w_{n}-2b$. If $d=\delta-a$, then $0<d<a$.
@@
We also record the following consequence of Lemma 3, which states a purely combinatorial fact. For any $v_{n}\in\mathcal{W}_{2}$, then
(8) $v_{n}-a=w_{n}-b-a\not\in\mathcal{W}_{0}\quad\text{for all }n.$
@@
Let us verify “reachability”. By Lemma 4, $v_{n}-b=w_{n}-2b\in\mathcal{W}_{0}\cup\mathcal{W}_{1}$. If $v_{n}-b\in\mathcal{W}_{1}$, then $(v_{n}-b)-a\in\mathcal{W}_{0}$ (by Ferguson pairing) and thus the option $v_{n}-b-a$ has value zero, by induction. Otherwise, if $v_{n}-b\in\mathcal{W}_{0}$ then by induction, the value is zero. Thus, it suffices to find an option to value one, given that $v_{n}-b\in\mathcal{W}_{0}$.
By (8), $v_{n}-a\notin\mathcal{W}_{0}$. Anti-collision of $\mathcal{W}_{2}$ rules out $v_{n}-a\in\mathcal{W}_{2}$. So $v_{n}$ has nim-value either one or three. If $v_{n}-a\in\mathcal{W}_{1}$ then it yields an option of value one by induction, and we are done. So, suppose position $v_{n}-a$ has value three. We claim that this implies that the option $v_{n}-a-b$ has value one. Firstly note that, by assumption and anti-collision, the position $(v_{n}-a)-b$ cannot have value three. And $v_{n}-(a+b)$ cannot have value two by induction and anti-collision of $\mathcal{W}_{2}$ (implied by Lemma 3). Moreover it cannot have value zero, by anti-collision, since this is the case where we assume that $v_{n}-b\in\mathcal{W}_{0}$. But then, since all other values are ruled out, the option $v_{n}-b-a$ must have value one. ∎
The second result of this section concerns the contribution of the set $\mathcal{W}_{3}:=(\mathcal{W}_{0}-\delta)\setminus\mathcal{W}_{0}$.
@@
Anti-collision is a consequence of Lemma 3, since $\mathcal{W}_{3}$ is a subset of an affine translation of $\mathcal{W}_{0}$.
For reachability, note that $\mathcal{W}_{0}-\delta=\mathcal{W}_{0}-b+a$, and hence, by subtracting $a$, a player can move to nim-value two. Thus, the set $\mathcal{W}_{3}$ cannot contain any nim-value two. Neither does it contain any nim-value zero (by definition) or one (by negated Ferguson pairing through $w_{0}-b\in\mathcal{W}_{2}$). ∎
It remains to show that $\mathcal{W}_{0}\sqcup\mathcal{W}_{1}\sqcup\mathcal{W}_{2}\sqcup\mathcal{W}_{3}=\mathbb{N}_{0}$, i.e. that these four sets exhaust all positions. To prove this, it suffices to show that the number of positions of $\mathcal{W}_{3}$ in one heap-period matches
(9) $a^{2}=(3\delta+a)a-3a\delta,$
the remaining positions in one heap-period. Let us explain. Obviously, the number of positions in a period of $\mathcal{W}_{0}$ or any translate of it is $a\delta$ (this is the index-period). From this, a direct calculation from the closed form $w_{n}$ shows that the period length of the $\mathscr{P}$-positions (the heap-period) is:
(10) $w_{n+a\delta}-w_{n}=a\delta+a\Bigl(\Bigl\lfloor\frac{n+a\delta}{a}\Bigr\rfloor-\Bigl\lfloor\frac{n}{a}\Bigr\rfloor\Bigr)+(a+\delta)\Bigl(\Bigl\lfloor\frac{n+a\delta}{\delta}\Bigr\rfloor-\Bigl\lfloor\frac{n}{\delta}\Bigr\rfloor\Bigr)=(3\delta+a)a.$
Note that, since $a\delta$ is a multiple of both $a$ and $\delta$, the two floor differences are $\delta$ and $a$, respectively.
The following section isolates the core number-theoretic mechanism.

### 29 Mar 2026 v2

@@ Abstract.
We determine the full nim-value structure of additive subtraction games in the primitive quadratic regime. The problem appears in Winning Ways by Berlekamp et al. in 1982; it includes a closed formula, involving Beatty-type bracket expressions on rational moduli, for determining the $\mathscr{P}$-positions, but to the best of our knowledge, a complete proof of this claim has not yet appeared in the literature; Miklós and Post (2024) established outcome-periodicity, but without reference to that closed formula. The primitive quadratic case captures the source of the quadratic complexity of the problem, a claim supported by recent research in the dual setting of sink subtraction by Bhagat et al. This study focuses on a number theoretic solution involving the classical closed formula, and we establish that each nim-value sequence resides on a linear shift of the classical $\mathscr{P}$-positions.
@@ 1. Introduction
Imagine a heap of counters sitting between two players, who take turns removing exactly $a$, $b$, or $a+b$ counters, where $0<a<b$ are given integers. The player who cannot move, because every legal removal would make the heap negative, loses. Simple to state, yet the question of who wins from any given heap size turns out to conceal surprisingly deep arithmetic.
This is an instance of a subtraction game [1], in the sub-class of additive subtraction. The tool for answering the winning question, not just for a single heap, but for any number of heaps played together, is the nim-value (or nimber) of a position. When two or more such games are played in parallel, with a player choosing on each turn which component game to move in, the combined game is called a disjunctive sum. See Figure 1.
Figure 1. A position of additive subtraction, with ruleset $S=\{2,5,7\}$, consisting of two heaps of sizes seven and eleven, played in a disjunctive sum. The player making the last move wins. Should Alice accept Bob’s offer to move first?
The famous Sprague–Grundy theory [3, 14, 10] tells us that every such component has a nim-value, a single non-negative integer, and the nim-value of the disjunctive sum is simply the bitwise XOR of the nim-values of its components. Suppose, for example, that you face two heaps, perhaps of different sizes, but both with nim-value three; the theory reveals that the position is losing for the player to move, independently of which heap they touch. The nim-values therefore reveal a complete winning strategy across any combination of rulesets and heap sizes.
The nim-value of a position is assigned recursively: it is the smallest non-negative integer not among the nim-values of all the move options; this is the classical mex (minimum excludant) rule. We think of it in terms of two components: for each position, one must verify both anti-collision (the same value does not appear among the options) and reachability (every smaller value appears among the options). For example, if a position can reach nim-values $\{0,1,3\}$, its own nim-value is two. Positions of nim-value zero are the $\mathscr{P}$-positions, the losing positions for the player to move; they are justified by anti-collision alone. For our three-move game $S=\{a,b,a+b\}$, the mex-rule implies all nim-values lie in $\{0,1,2,3\}$, since at most three options are available from any position. The question is: “which positions get which value?”
The recursive algorithm answers this for small heap sizes. Let us illustrate it for the game in Figure 1, computing just enough to guide Alice towards the right answer to Bob’s question.
Table 1. Nim-values for the ruleset $S=\{2,5,7\}$.
Heap size 0 1 2 3 4 5 6 7 8 9 10 11
Nim-value 0 0 1 1 0 2 1 3 2 2 0 3
Both heaps have nim-value three, so their XOR is $3\oplus 3=0$: this is a $\mathscr{P}$-position, and Alice should decline Bob’s offer. But for large heaps, or many heaps, direct computation becomes infeasible.
We prove here that each nimber sequence has an elegant closed form formula, which makes computation tractable. [^1] Since it is known that all subtraction games are periodic [7], one could argue that every subtraction game is tractable, even without knowledge of a closed formula. However, Flammenkamp [4] conjectures that there exist sequences of games with exponential period length in terms of $\max S$, which, if true, would slow down any such semi-tractable approach considerably. The sequence of $\mathscr{P}$-positions has the following form in the general case including $S=\{2,5,7\}$. For a given instance of additive subtraction with game parameters $0<a<b$, for all $n\in\mathbb{N}=\{0,1,\ldots\}$, let
(1) $w_{n}:=n+a\Bigl\lfloor\frac{n}{a}\Bigr\rfloor+b\Bigl\lfloor\frac{n}{\delta}\Bigr\rfloor,$
where $\delta=b-a$. A slight generalization of this formula, involving two interlocking floor functions, first appeared in [1] in 1982, but without a proof. To the best of our knowledge, no proof has appeared in the literature since, although it appeared in a similar form in [4] in the late 1990s. The present paper closes that gap and goes further: we determine the complete nim-value structure. Note that the nim-value of any position smaller than $a$ is zero, which also follows directly by (1), since both floor terms evaluate to zero.
We work in the primitive quadratic regime, $a<\delta<2a$ with $\gcd(a,\delta)=1$, so we study the ruleset family primitive quadratic additive subtraction. This is the first parameter range in which the two scales $a$ and $\delta$ interact non-trivially, and it captures the core complexity of the problem; see also the dual setting of Bhagat et al. [2]. We use the following notation throughout. If $X\subset\mathbb{N}$ and $y\in\mathbb{N}$, then
$X+y:=\{x+y:x\in X\}\quad\text{and}\quad X-y:=\{x-y:x\in X,\,x\geqslant y\},$
so that XyX-y contains only non-negative integers. Writing $\mathcal{W}_{0}=\{w_{n}:n\geqslant 0\}$, the four nim-value classes are
$\mathcal{W}_{0},\quad\mathcal{W}_{1}=\mathcal{W}_{0}+a,\quad\mathcal{W}_{2}=\mathcal{W}_{0}-b,\quad\mathcal{W}_{3}=(\mathcal{W}_{0}-\delta)\setminus\mathcal{W}_{0}.$
In words: three of the four classes are simple shifts of the $\mathscr{P}$-positions, while the fourth, the nim-value three positions, and the most intricate, is what remains after removing the $\delta$-shift’s overlap with $\mathcal{W}_{0}$.
@@
Assume $a<\delta<2a$ and $\gcd(a,\delta)=1$, with $b=a+\delta$. Then $\mathcal{W}_{0},\mathcal{W}_{1},\mathcal{W}_{2},\mathcal{W}_{3}$ partition $\mathbb{N}$, and $\mathcal{W}_{i}$ is precisely the set of positions of nim-value $i$, for $i\in\{0,1,2,3\}$.
Three of the four classes are handled with relatively fewer complications: $\mathcal{W}_{0}$ by a number-theoretic anti-collision argument, $\mathcal{W}_{1}$ by Ferguson’s pairing principle [1] (the smallest move $a$ pairs every $\mathscr{P}$-position with a unique nim-value one position), and $\mathcal{W}_{2}$ by a straightforward induction on reachability combined with nim-value exclusion. The nim-value three positions in $\mathcal{W}_{3}$ are the most delicate: we must establish their count per period, and we use a collision-counting argument that forms the arithmetic heart of the paper.
We call (1) a “bracket expression on three terms”, and more generally, an expression
$\sum_{i=1}^{t}c_{i}\lfloor n\alpha_{i}\rfloor,$
with integer coefficients $c_{i}$ and positive rational moduli $\alpha_{i}$, is a bracket expression on $t$ terms.
As we will see, the arithmetic structure of the problem ultimately stems from the index-regular gap structure of the $\mathscr{P}$-position sequence (1): only four gap sizes appear, and their locations are determined exactly by the residue classes of the index modulo $a$ and $\delta$. The index-period length of (1) is $a\delta$, which implies that the heap-period length is quadratic in $a$, namely $h:=(3\delta+a)a$. The densities of the four classes are therefore
$\frac{\delta}{3\delta+a},\quad\frac{\delta}{3\delta+a},\quad\frac{\delta}{3\delta+a},\quad\frac{a}{3\delta+a},$
for $\mathcal{W}_{0},\mathcal{W}_{1},\mathcal{W}_{2},\mathcal{W}_{3}$ respectively. Notably, the nim-value three positions are not only the most intricate, they are also the rarest, with density strictly less than the other three classes, reflecting the primitive quadratic inequality $a<\delta$.
Outside the primitive quadratic regime, essentially nothing new occurs: either one denominator becomes a multiple of the other in (1), which reduces the problem to linear complexity, or certain pairs of nim-value “factors” expand regularly by $2a$ with each increase of $b$ by $2\delta_{0}$ (if $\delta_{0}$ is defined by the primitive quadratic regime), preserving quadratic complexity; see [2].
@@
The paper is organized as follows. Section 2 establishes anti-collision for the candidate set of $\mathscr{P}$-positions, $\mathcal{W}_{0}$, proving that no move connects any two elements of $\mathcal{W}_{0}$, and deduces the nim-value one structure of $\mathcal{W}_{1}$, via Ferguson’s pairing principle. Section 3 handles the nim-value two and three positions, with the key reachability arguments supported by Figure 2, which gives a geometric interpretation of the central arithmetic mechanism. Section 4 isolates the number-theoretic heart of the paper via a collision-counting argument. We demonstrate that, in one heap-period, exactly $ad$ elements of $\mathcal{W}_{0}$ lie within distance $\delta$ of another element of $\mathcal{W}_{0}$, from which the nim-value three count follows. Section 5 illuminates the number-theoretic approach with three explicit numerical examples. Section 6 assembles all ingredients into the proof of Theorem 1, and Section 7 closes with open problems and directions for future work.
@@ 2. Anti-collision of candidate $\mathscr{P}$-positions
In this section, we prove the first part of the mex-verification, anti-collision, for the candidate set of $\mathscr{P}$-positions. Note that the reachability part is empty, since we here study zero values. However, since any ‘independent subset of the natural numbers’ satisfies the anti-collision requirement, the true positioning of the zero values will only be confirmed when we have established the complete nim-value tetra-partitioning of all game positions, as in Theorem 1.
@@
Let $S=\{a,b,a+b\}$ with $b=a+\delta$ and $a<\delta<2a$. For all $n\in\mathbb{N}$, let
@@
If $\beta\geqslant 1$ then $b(\beta-1)\geqslant 0$, and since $a\alpha\geqslant 0$ and $\ell\geqslant 1$, we reach a contradiction. Hence $\beta=0$, which gives
@@
Recall the affine translation of $\mathcal{W}_{0}$, the position set $\mathcal{W}_{1}=\{w_{n}+a:n\geqslant 0\}$. By Lemma 3, $\mathcal{W}_{1}$ satisfies the anti-collision property. Moreover, every position $x$ in this set has an option in $\mathcal{W}_{0}$, namely xax-a. If the set $\mathcal{W}_{0}$ constitutes the set of $\mathscr{P}$-positions, then, since $a=\min S$, Ferguson’s pairing principle [1] implies that every position in $\mathcal{W}_{1}$ has nim-value one, and there is no other position of value one. Indeed, by induction, we can already deduce that each $x\in\mathcal{W}_{1}$ has nim-value one. (The other direction will follow only as we prove the main theorem.)
@@ 3. Higher nim-values
In this section, we study positions with nim-values two and three. For all $n\geqslant 0$, let $v_{n}:=w_{n+\delta}-b$. Then $\mathcal{W}_{2}:=\{v_{n}:n\geqslant 0\}$.
The following lemma identifies the key arithmetic property of positions in $\mathcal{W}_{2}$. It makes explicit the effect of subtracting $b$ from any element of $\mathcal{W}_{2}$. Observe that, since $w_{\delta}=2b$ and by monotonicity,
(7) $v_{n}-b=w_{n+\delta}-2b\geqslant 0.$
If $d=\delta-a$, then $0<d<a$.
@@
We also record the following consequence of Lemma 3, which states a purely combinatorial fact. For any $v_{n}\in\mathcal{W}_{2}$, then, for all $n\geqslant 0$, by anti-collision in the set $\mathcal{W}_{0}$,
(8) $v_{n}-a=w_{n+\delta}-b-a\not\in\mathcal{W}_{0}.$
Recall the set $\mathcal{W}_{3}:=(\mathcal{W}_{0}-\delta)\setminus\mathcal{W}_{0}$. We will establish that every position in $\mathcal{W}_{3}$ has nim-value three; First we establish that all three options are available (as required for a position of nim-value three).
Lemma 5.
The following hold:
  • (i)
    If $n<a$, then $w_{n}<\delta$.
  • (ii)
    If $a\leqslant n<\delta$, then $w_{n}-\delta\in\mathcal{W}_{0}$.
  • (iii)
    If $n\geqslant\delta$, then $w_{n}-\delta\geqslant a+b$.
In particular, every element $x\in\mathcal{W}_{3}$ satisfies $x\geqslant a+b$, so all three moves are available from any position in $\mathcal{W}_{3}$.
Proof.
Item (i): if $n<a$ then $\lfloor n/a\rfloor=\lfloor n/\delta\rfloor=0$, so $w_{n}=n<a<\delta$.
Item (ii): if $a\leqslant n<\delta$ then $\lfloor n/\delta\rfloor=0$ and $\lfloor n/a\rfloor=1$, so $w_{n}=n+a$ and $w_{n}-\delta=n-d$. Since $n<\delta=a+d$, we have nd<an-d<a, and hence $\lfloor(n-d)/a\rfloor=\lfloor(n-d)/\delta\rfloor=0$, giving $w_{n-d}=n-d$. Thus $w_{n}-\delta=w_{n-d}\in\mathcal{W}_{0}$.
Item (iii): if $n\geqslant\delta$ then $\lfloor\delta/a\rfloor=\lfloor\delta/\delta\rfloor=1$, so $w_{n}\geqslant w_{\delta}=\delta+a+b$, by $(w_{n})$ strictly increasing. Thus, $w_{n}-\delta\geqslant a+b$ for all $n\geqslant\delta$.
The final claim follows from (i) and (ii): no element of $W_{3}$ can arise from $n<a$ (as $w_{n}-\delta<0$) or from $a\leqslant n<\delta$ (as $w_{n}-\delta\in W_{0}$), so every element of $W_{3}$ has $n\geqslant\delta$, and by (iii) satisfies $x\geqslant a+b$. ∎
@@
Let us verify “reachability”. By Lemma 4 and (7), $v_{n}-b=w_{n+\delta}-2b\in\mathcal{W}_{0}\cup\mathcal{W}_{1}$. If $v_{n}-b\in\mathcal{W}_{1}$, then, by definition of $\mathcal{W}_{1},v_{n}-b\geqslant a$, and thus, by Ferguson pairing, $(v_{n}-b)-a\in\mathcal{W}_{0}$, with value zero, by induction. Otherwise, if $v_{n}-b\in\mathcal{W}_{0}$ then by induction, the value is zero.
Thus, it suffices to find an option of $v_{n}$ to value one, given that
(9) $v_{n}-b\in\mathcal{W}_{0}.$
Let us study the option $v_{n}-a$. By (8), $v_{n}-a\notin\mathcal{W}_{0}$. Anti-collision of $\mathcal{W}_{2}$, by induction, implies $v_{n}-a\not\in\mathcal{W}_{2}$. So, by induction, $v_{n}-a$ has nim-value either one or three. If $v_{n}-a\in\mathcal{W}_{1}$ then it yields an option of value one by induction, and we are done. So, suppose position
(10) $v_{n}-a\in\mathcal{W}_{3}.$
We claim that this implies that there is an option $v_{n}-a-b\geqslant 0$ of value one (by induction). The bounding away from zero is clear by Lemma 5. Moreover, note that, by the assumption (10), induction and anti-collision, the position $(v_{n}-a)-b$ cannot have value three. And $v_{n}-(a+b)$ cannot have value two by induction and anti-collision of $\mathcal{W}_{2}$ (implied by Lemma 3). Moreover it cannot have value zero, by anti-collision, since this is the case (9), where we assume that $v_{n}-b\in\mathcal{W}_{0}$. But then, since all other values are ruled out, the option $v_{n}-b-a$ must have value one. ∎
The second result of this section concerns the contribution of the set $\mathcal{W}_{3}$. We will prove that all those positions have nim-value three. Recall that, by our convention, $\mathcal{W}_{0}-\delta\subset\mathbb{N}$. Note that $w_{n}-\delta\geqslant 0$ if and only if $n\geqslant a$. Let us begin with a technical reachability lemma, which concerns a special class of positions in $\mathcal{W}_{3}$. In Figure 2, we have depicted the main idea of the proof.
Lemma 7.
Let $n=qa+r\geqslant a$ with $0\leqslant r<a$, and consider the position $w_{n}-\delta$. Given the provisos $r<d:=\delta-a$ and $w_{n}-\delta\notin\mathcal{W}_{0}$, the option $(w_{n}-\delta)-b\in\mathcal{W}_{0}$.
Proof.
We set out to prove that, given the provisos in the statement, then $w_{n}-w_{n-d}=2b-a$, which implies the lemma.
The proof is by way of contradiction, and to this purpose we define another representation of the index $n=p\delta+s$, with $0\leqslant s<\delta$. We first demonstrate that, given the two provisos, then $s<d$. Suppose for contradiction that $s\geqslant d$. Since $n-d=p\delta+s-d$, then no multiple of $\delta$ lies in the interval (nd,n](n-d,n], namely $p\delta\leqslant n-d\leqslant n=p\delta+s<(p+1)\delta$. Hence
(11) $\left\lfloor\frac{n}{\delta}\right\rfloor-\left\lfloor\frac{n-d}{\delta}\right\rfloor=0.$
Since $r<d<a$, the interval (nd,n](n-d,n] contains exactly one multiple of $a$, namely $qa$; this follows since $n-d=qa+r-d<qa\leqslant n=qa+r$, as $r<d$. Thus
(12) $\left\lfloor\frac{n}{a}\right\rfloor-\left\lfloor\frac{n-d}{a}\right\rfloor=1.$
Altogether (11) and (12) give
$w_{n}-w_{n-d}$ $=d+a\cdot 1+b\cdot 0$
$=d+a$
$=\delta,$
and hence $w_{n}-\delta=w_{n-d}\in\mathcal{W}_{0}$, contradicting the proviso $w_{n}-\delta\notin\mathcal{W}_{0}$. Thus $s<d$.
Since $s<d$, then $s+a<\delta$. Thus,
$\lfloor(n-d)/\delta\rfloor$ $=\lfloor(q\delta+s-\delta+a)/\delta\rfloor$
$=\lfloor(q\delta+s+a)/\delta\rfloor-1$
$=\lfloor n/\delta\rfloor-1,$
and hence $\lfloor n/\delta\rfloor-\lfloor(n-d)/\delta\rfloor=1$. This will contribute the term $b$ to $w_{n}-w_{n-d}$, and by (12), $\lfloor n/a\rfloor-\lfloor(n-d)/a\rfloor=1$, which contributes the amount $a$. Altogether, we get
$w_{n}-w_{n-d}$ $=d+a\cdot 1+b\cdot 1$
$=\delta+b$
$=2b-a,$
and hence $w_{n-d}=w_{n}-2b+a\in\mathcal{W}_{0}$. ∎
[…]
Figure 2. The Winning Model. The picture requires $w_{n}-\delta\not\in\mathcal{W}_{0}$. Zero is at 12 o’clock modulo $a$ or $\delta$ respectively; the wheels have $a=7$ and $\delta=12$ equally spaced ticks, respectively. The colored arcs span $d=5$ units on each wheel, representing the zones of sufficently small residues. The red dots show residues $r=2$ and $s=4$, both satisfying $r,s<d$. Since both residues jump across the north pole when the index shifts by $d$, the bracket expression arithmetic produces a combined heap change of $d+a+b=\delta+b$, which suffices to reach a $\mathscr{P}$-position by subtracting $b$ from $w_{n}-\delta$.
@@
Recall that $\mathcal{W}_{3}:=(\mathcal{W}_{0}-\delta)\setminus\mathcal{W}_{0}$. Anti-collision holds since $\mathcal{W}_{3}$ is a subset of a linear shift of $\mathcal{W}_{0}$. Further, note that $\mathcal{W}_{0}-\delta=\mathcal{W}_{0}-b+a$, and hence, by subtracting $a$ from a position $x\geqslant a$ in $\mathcal{W}_{3}$, by induction, a player can move to nim-value two. Let us consider reachability to the smaller nim-values. Let $n\equiv r\pmod{a}$, and let $d=\delta-a$. By observing also Lemma 5, we record the following claims:
  • (i)
    If $w_{n}-\delta\in\mathcal{W}_{3}$ with $r<d$ then, subtracting $b$ gives value zero, while subtracting $(a+b)$ gives value one.
  • (ii)
    If $w_{n}-\delta\in\mathcal{W}_{3}$ with $r\geqslant d$ then, subtracting $b$ gives value one, while subtracting $(a+b)$ gives value zero.
In case (i), we use Lemma 4 to see that the value one option is correct, by induction. Indeed, $w_{n}-b+a-(b+a)=w_{n}-2b=w_{n-\delta}+a\in\mathcal{W}_{1}$. Therefore, it suffices to prove that, given the provisos $r<d$ and $w_{n}-\delta=w_{n}-b+a\not\in\mathcal{W}_{0}$, the position
(13) $w_{n}-2b+a=w_{n-d},$
which by induction implies value zero. By Lemma 7, equation (13) holds, and we are done with this case.
In case (ii), we reach value zero by subtracting $a+b$ from $w_{n}-\delta$. Indeed, by invoking the proviso $r\geqslant d$, Lemma 4 gives that $w_{n}-2b=w_{n-\delta}\in\mathcal{W}_{0}$ has value zero, by induction. Moreover, by instead subtracting $b$, Ferguson pairing gives value one at heap size $w_{n}-2b+a=w_{n-\delta}+a\in\mathcal{W}_{1}$, by induction, and since $a=\min S$. ∎
We now understand the nim-values of the relevant sets. It remains to show that, in fact, $\mathcal{W}_{0}\sqcup\mathcal{W}_{1}\sqcup\mathcal{W}_{2}\sqcup\mathcal{W}_{3}=\mathbb{N}$, i.e. these four sets exhaust all positions. To prove this, it suffices to show that the number of positions of $\mathcal{W}_{3}$, in one heap-period, matches
(14) $a^{2}=(3\delta+a)a-3a\delta.$
These are the remaining positions in one heap-period, given that the three sets $\mathcal{W}_{0},\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ have been accounted for. Let us explain. Obviously, the number of positions in a period of $\mathcal{W}_{0}$, or any translate of it, is $a\delta$ (this is the index-period). From this, a direct calculation from the closed form $w_{n}$ shows that the period length of the $\mathscr{P}$-positions (the heap-period) is:
$w_{n+a\delta}-w_{n}$ $=a\delta+a\Bigl(\Bigl\lfloor\frac{n+a\delta}{a}\Bigr\rfloor-\Bigl\lfloor\frac{n}{a}\Bigr\rfloor\Bigr)+(a+\delta)\Bigl(\Bigl\lfloor\frac{n+a\delta}{\delta}\Bigr\rfloor-\Bigl\lfloor\frac{n}{\delta}\Bigr\rfloor\Bigr)$
(15) $=(3\delta+a)a.$
Note that, since $a\delta$ is a multiple of both $a$ and $\delta$, the two floor differences are $\delta$ and $a$, respectively. The following section isolates the core number-theoretic mechanism in establishing the count in (14).